# Distribution is Unlimited. PA# AFRL-2024-2340

### v1

@@ III.2 Generalizing with Weak Pairings
$\mu(A)=\max_{\norm{v}=1}\langle\!\langle Av,v\rangle\!\rangle$ (12)
We will not discuss weak pairings nor lognorms in great depth, only the most relevant properties. The basic properties of lognorms and weak pairings are defined in Appendix C and Appendix D respectively. More detailed treatment can be found in the references therein. Broadly speaking, the lognorms $\mu(A)$ will be used to provide bounds for the weak pairings as well as feasible methods of calculation. These weak pairings allow for the definition of a family of candidate Lyapunov functions on the VDS of (7), indexed by $\ell\in\mathbb{R}^{\geq 1}$, as $\{V_{\ell}(z):V_{\ell}(z)=\frac{1}{2}\langle\!\langle z,z\rangle\!\rangle_{\ell}\geq 0\}$ with the respective evolution $\dot{V}_{\ell}(\dot{z},z)=\langle\!\langle\dot{z},z\rangle\!\rangle_{\ell}$. If any of these weak pairings yields $\dot{V}_{\ell}(\dot{z},z)\leq 0$ for all $z$ and $u$, then that $V_{\ell}(z)$ is a Lyapunov function and the VDS is stable and GS occurs. Additionally, if $\frac{\dot{V}_{\ell}(z)}{V_{\ell}(z)}<0$ for all non-zero $z$ and $\dot{V}_{\ell}(0)=0$, then the system is globally exponentially stable with a unique equilibrium at $z=0$. The stability condition for the proposed family of Lyapunov functions has the following form, which guarantees global exponential stability.
@@
$\qquad+\frac{\langle\!\langle\sigma(Ax+Bu)-\sigma(Ay+Bu),x-y\rangle\!\rangle}{\norm{x-y}^{2}}<0$
The first weak pairing above can be bounded by $\mu(-C)$ by definition. Using the shorthand $\sigma(Ax+Bu)=\hat{\sigma}(x)$ for the activation function and substitution $x\to y+hv$ for conciseness, the second weak pairing term of (13) can be bounded by the maximal lognorm over all possible Jacobians $D\hat{\sigma}(x)$ of the activation function.
@@
$=\max_{x}\mu(D\hat{\sigma}(x))$
The sufficient condition for global exponential stability of the VDS is $\max_{x}\mu(D\hat{\sigma}(x))<-\mu(-C)$. The calculation of $\max_{x}\mu(D\hat{\sigma}(x))<-\mu(-C)$ is difficult for the arbitrary norms, but the special cases of the 1-lognorm and the $\infty$-lognorm (that is, $\ell=1$ and $\ell=\infty$) with saturating activation function $\sigma$ (such as tanh), there is a closed form solution in terms of the connectivity matrix $A$ [28]. The maximum of these lognorms over all possible Jacobians $D\hat{\sigma}(x)$ yield the following analytic form, with the usage of the entry-wise or Hadamard product $(I\circ A)_{ij}=(I)_{ij}(A)_{ij}$. The proofs and details of this are in Appendix E.
$\max_{x}\mu_{1}(D\hat{\sigma}(x))=\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}$ (15)
$\max_{x}\mu_{\infty}(D\hat{\sigma}(x))=\max\{\mu_{\infty}(A),0\}$
It suffices that imposing either $\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}<-\mu(-C)$ or $\max\{\mu_{\infty}(A),0\}<-\mu(-C)$ is enough to guarantee the ESP and GS. These conditions may be readily calculated according to the definitions of $\mu_{1}(A)$ and $\mu_{\infty}(A)$ in (17) and (20) respectively. These lognorms provide a tighter bound than their matrix norm counterparts $\norm{A}_{1}$ and $\norm{A}_{\infty}$ respectively, as shown at the end of Appendix E. If $C$ is chosen to be the identity, these conditions simplify and are given by the following.
$\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}<1$ (16)
$\mu_{\infty}(A)<1$
Only the 1-lognorm and $\infty$-lognorm cases are checked in the approach detailed here. When neither of these conditions are satisfied, it is still possible that contraction occurs through some other lognorm, but it is not yet known how to acquire those sufficient conditions. From this analysis, it is clear that the leak term $C$ must be present for contraction to be guaranteed with logarithmic norms. This is due to both $\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}$ and $\max\{\mu_{\infty}(A),0\}$ always being non-negative. The stability provided by the leak term cannot be overlooked if ESP and GS are to be guaranteed with this approach. Nevertheless, these are only sufficient conditions and it is still possible for such systems to show ESP and GS without satisfying these conditions.
@@ III.3 Interpretation
The first condition, corresponding to the 1-lognorm in (16), can be further broken down into two parts – $\mu_{1}(A)<1$ and $\mu_{1}(A-I\circ A)<1$. Both sub-conditions must be true in order for the condition to hold. The $1$-lognorm of an arbitrary matrix $M$ can be calculated via the following.
$\mu_{1}(M)=\max_{j}\left(m_{jj}+\sum_{i\neq j}\absolutevalue{m_{ij}}\right)$ (17)
For the connectivity matrix $A$, $\mu_{1}(A)<1$ can be easily calculated according to (17), where the column-sum is over the the magnitude of the off-diagonal terms. Effectively, this imposes that for every column (indexed by $j$), the sum the off-diagonal |$\absolutevalue{m_{ij}}$ | must be less than $1-a_{ij}$.
$\sum_{i\neq j}\absolutevalue{a_{ij}}<1-a_{jj}$ (18)
This is consistent with the common practice of choosing $A$ to be sparse, since the $\sum_{i\neq j}\absolutevalue{a_{ij}}$ | term grows quickly with a denser $A$. This condition suggests that a more-negative diagonal element is desirable, which ties in the concept of dissipation and its role synchronization ((NEED CITE)). In addition, this condition fails to hold when $a_{ij}\geq 1$, due to the sum of absolute values always being non-negative.
The second sub-condition $\mu_{1}(A-I\circ A)<1$ appears more complicated, but is conceptually simple as well. Let the Hadamard product be represented by the operator $\circ$, so $I\circ A$ constructs a diagonal matrix with the diagonal elements of $A$. The resulting matrix $A-I\circ A$ is similar to $A$, but with its diagonal elements set to zero. This sub-condition is a special case of (18), given by imposing the following constraint for every column.
$\sum_{i\neq j}\absolutevalue{a_{ij}}<1$ (19)
Though (18) and (19) appear to overlap, these two sub-conditions disagree when the diagonal elements $m_{jj}$ are positive. Otherwise, if $m_{jj}$ are all non-positive, then (19) implies (18).
The second condition for stability is $\max\{\mu_{\infty}(A),0\}<1$. Zero appears in this equation only to indicate that the leak term with matrix $C=I$, discussed in Section III.2, must be present to guarantee ESP and GS. Adhering to this constraint, the condition is effectively $\mu_{\infty}(A)<1$. This corresponds to the $\infty$-lognorm, which can be calculated by the following.
$\mu_{\infty}(M)=\max_{i}\left(m_{ii}+\sum_{j\neq i}\absolutevalue{m_{ij}}\right)$ (20)
The above condition is the row-wise version of (17), giving the corresponding restriction (18) when applied to the connectivity matrix $A$.
$\sum_{j\neq i}\absolutevalue{a_{ij}}<1-a_{ii}$ (21)
@@ IV.1 UAP in Discrete Time
Take the discrete time reservoir outlined in (1), we have established that $A:\norm{A}<1$ and $B$ can be arbitrarily chosen such that the ESP holds. The generation of trajectories is often called the listening phase and implies the existence of a unique time-invariant function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$ such that $x_{t}=\Phi(u_{t})$ as $t\to\infty$, but without an explicit representation of $\Phi$.
$x_{t+1}$ $=\sigma(Ax_{t}+Bu_{t})$ (22)
$z_{t}$ $=Wx_{t}+b$
@@
The constraint $\norm{A}<1$ allows for only one basin of attraction in the reservoir, resulting in the map $\Phi$ induced by the dynamics to also be unique. Aside from some transients, $u$ can be recovered from $x$ through the application of $\Phi^{-1}$, approximated by $W$. Further suppose that the approximation of $\Phi^{-1}$ by $W$ holds arbitrarily well, such that $u_{T}=\Phi^{-1}(x_{T})=Wx_{T}$ for a finite $T$. As such, the second line of (22) can be substituted into the first, resulting in an autonomous reservoir that is used for $t>T$. With the shorthand $\bar{A}=A+BW$ and $\bar{b}=Bb$, this newly formulated and autonomous reservoir is usually called the predicting phase.
@@
$=\sigma(\bar{A}x_{t}+\bar{b})$
$\hat{u}_{t}$ $=Wx_{t}+b$
It is surprising that a non-autonomous system like (22) can be ‘converted’ into an autonomous system with a linear readout in the manner described in (23). Often overlooked is that the form of (23) admits the UAP in the feedforward sense of Cybenko [30], Hornik et al. [31], or Leshno et al. [32]. Approximation theorems for recurrent networks are not necessary.
Also overlooked is the relationship between contracting functions and the activation functions compatible with the UAP. Nonlinear contracting functions used as activation functions will always be compatible with the UAP. This is most succinctly shown with the results of Leshno et al. [32] which states that the UAP holds in networks of the form (23) if and only if the activation function is nonlinear and non-polynomial.
The proof of nonlinear contracting functions being non-polynomials is outlined as follows. If two distinct points on a continuously differentiable function cannot generate a slope greater than unity, then this function is globally contracting. Degree one and zero polynomials are linear, so they are automatically excluded from the set of nonlinear contracting functions. Polynomial functions of degree two or higher are nonlinear but will always manifest an unbounded slope eventually, so they cannot be contracting. Thus, the set of nonlinear contracting functions is mutually exclusive with the set of nonlinear polynomial functions. Therefore, all nonlinear contracting functions are non-polynomials.
@@
The only invoked property of the activation function $\sigma$ thus far was nonlinear global contraction, which is a special case of Lipschitz continuous. Other commonly invoked properties of sigmoid functions – monotonicity and saturation – are not used in this work, nor in Cybenko [30]. The UAP requires a large but finite number of nodes, hence the condition that $N\gg M$. The embedding criteria of Whitney (consequently of Takens as well) whereby $\dim N>2\dim M$ is almost surely satisfied in the large $N$ limit. This large $N$ limit also contributes to the quality of the approximation $\Phi^{-1}(x_{t})\simeq Wx_{t}+b$ for ergodic systems [33].
@@ IV.2 UAP in continuous time
For continuous time reservoirs, let us first consider the Hopfield-like equations outlined in (7) with a leak current. We constrain $A$ such that $\norm{A}<1$ and $C=I$, but $B$ can be arbitrarily chosen such that the ESP holds. This implies the existence of some time-invariant function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$ where $x(t)=\Phi(u(t))$ as $t\to\infty$ without explicit knowledge about the representation of $\Phi$.
$\dot{x}$ $=-x+\sigma(Ax+Bu)$ (24)
$\hat{u}$ $=Wx+b$
Much like the discrete time case, trajectories are generated using the above dynamics, but an arbitrary time integration scheme is necessary. The listening and training phases are conducted as usual, where the weights $W$ and biases $b$ are solved for. The prediction phase is also done by ‘converting’ the non-autonomous system into an autonomous one, via the closing of the input-output loop.
$\dot{x}$ $=-x+\sigma(\bar{A}x+\bar{b})$ (25)
$\hat{u}$ $=Wx+b$
@@
$\dot{x}$ $=\sigma(\bar{A}x+\bar{b})$ (26)
$\hat{u}$ $=Wx+b$
Matrix $A$ cannot yet be chosen such that the ESP is a priori guaranteed with the contraction approach following the arguments at the end of Section 3. However, $A$ can be chosen such that the ESP is observed empirically for some input $u$ of interest. Assuming so, the reservoir trajectories are generated and the weights $W$ and bias $b$ are acquired by training. The resulting dynamics of the autonomous reservoir can be interpreted as a NODE with a linear output layer, but introduces another constraint in that the activation function are also invertible. There are multiple results [36, 37, 38, 39] that have established narrower forms of UAP for NODEs, which are inherited by autonomous reservoirs (26) with invertible activation functions. In particular, Zhang et al. [36] show that some augmented NODEs – augmented by imposing that the network has number of neurons at least twice the number of inputs – are capable of universally approximating homeomorphism of a topological space onto itself. Teshima et al. [37] show that NODEs are universal approximators of diffeomorphisms of a topological space onto itself.

### v2

@@ III.2 Generalizing with Weak Pairings
$\mu(A)=\max_{\norm{v}=1}\langle\!\langle Av,v\rangle\!\rangle.$ (12)
We will not discuss weak pairings nor lognorms in great depth, only the most relevant properties. The basic properties of lognorms and weak pairings are defined in Appendixes C and D respectively. More detailed treatment can be found in the references therein.
Broadly speaking, the lognorms $\mu(A)$ will be used to provide bounds for the weak pairings as well as feasible methods of calculation. These weak pairings allow for the definition of a family of candidate Lyapunov functions on the VDS of (7), indexed by $\ell\in\mathbb{R}^{\geq 1}$, as $\{V_{\ell}(z):V_{\ell}(z)=\frac{1}{2}\langle\!\langle z,z\rangle\!\rangle_{\ell}\geq 0\}$ with the respective evolution $\dot{V}_{\ell}(\dot{z},z)=\langle\!\langle\dot{z},z\rangle\!\rangle_{\ell}$. If any of these weak pairings yields $\dot{V}_{\ell}(\dot{z},z)\leq 0$ for all $z$ and $u$, then that $V_{\ell}(z)$ is a Lyapunov function and the VDS is stable and GS occurs.
Additionally, if $\dot{V}_{\ell}(0)=0$ and $\frac{\dot{V}_{\ell}(z)}{V_{\ell}(z)}<0$ for all nonzero $z$, then the system is globally exponentially stable with a unique equilibrium at $z=0$. The stability condition for the proposed family of Lyapunov functions has the following form, which guarantees global exponential stability:
@@
$\qquad+\frac{\langle\!\langle\sigma(Ax+Bu)-\sigma(Ay+Bu),x-y\rangle\!\rangle}{\norm{x-y}^{2}}<0.$
The first weak pairing term above can be bounded by $\mu(-C)$ by definition. Using the shorthand $\sigma(Ax+Bu)=\hat{\sigma}(x)$ for the activation function and substitution $x\to y+hv$ for conciseness, the second weak pairing term of (13) can be bounded by the maximal lognorm over all possible Jacobians $D\hat{\sigma}(x)$ of the activation function:
@@
$=\max_{x}\mu(D\hat{\sigma}(x)).$
The sufficient condition for global exponential stability of the VDS is $\max_{x}\mu(D\hat{\sigma}(x))<-\mu(-C)$. The calculation of this quantity is difficult for arbitrary norms, except for the special cases of the 1-lognorm and the $\infty$-lognorm (that is, $\ell=1$ and $\ell=\infty$) with saturating activation function $\sigma$ (such as tanh). These case admit a closed form solution in terms of the connectivity matrix $A$ [29]. The maximum of these lognorms over all possible Jacobians $D\hat{\sigma}(x)$ yields the following analytic form, with the usage of the entry-wise or Hadamard product $(I\circ A)_{ij}=(I)_{ij}(A)_{ij}$. The proofs and details of this are in Appendix E:
$\max_{x}\mu_{1}(D\hat{\sigma}(x))=\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\},$ (15)
$\max_{x}\mu_{\infty}(D\hat{\sigma}(x))=\max\{\mu_{\infty}(A),0\}.$
It suffices that imposing either $\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}<-\mu(-C)$ or $\max\{\mu_{\infty}(A),0\}<-\mu(-C)$ is enough to guarantee the ESP and GS. These conditions may be readily calculated according to the definitions of $\mu_{1}(A)$ and $\mu_{\infty}(A)$ in (17) and (20) respectively. These lognorms provide a tighter bound than their matrix norm counterparts $\norm{A}_{1}$ and $\norm{A}_{\infty}$ respectively, as shown at the end of Appendix E. If $C$ is chosen to be the identity, these conditions simplify and are given by the following:
$\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}<1,$ (16)
$\mu_{\infty}(A)<1.$
Only the 1-lognorm and $\infty$-lognorm cases are checked in the approach detailed here. When neither of these conditions are satisfied, it is still possible that contraction occurs through some other lognorm, but it is not yet known how to acquire those sufficient conditions. From this analysis, it is clear that the leak term $C$ must be present for contraction to be guaranteed with logarithmic norms. This is due to both $\max\{\mu_{1}(A),\mu_{1}(A-I\circ A)\}$ and $\max\{\mu_{\infty}(A),0\}$ always being nonnegative. The stability provided by the leak term cannot be overlooked if ESP and GS are to be guaranteed with this approach. Nevertheless, these are only sufficient conditions and it is still possible for reservoir systems to show ESP and GS without satisfying these conditions.
@@ III.3 Interpretation
The first condition, corresponding to the 1-lognorm in (16), can be further broken down into two parts – $\mu_{1}(A)<1$ and $\mu_{1}(A-I\circ A)<1$. Both subconditions must be true for the condition to hold. The $1$-lognorm of an arbitrary matrix $M$ can be calculated via the following:
$\mu_{1}(M)=\max_{j}\left(m_{jj}+\sum_{i\neq j}\absolutevalue{m_{ij}}\right).$ (17)
For the connectivity matrix $A$, $\mu_{1}(A)<1$ can be easily calculated according to (17), where the column-sum is over the the magnitude of the off-diagonal terms. Effectively, this imposes that for every column (indexed by $j$), the sum the off-diagonal |$\absolutevalue{m_{ij}}$ | must be less than $1-a_{ij}$:
$\sum_{i\neq j}\absolutevalue{a_{ij}}<1-a_{jj}.$ (18)
This is consistent with the common practice of choosing $A$ to be sparse, since the $\sum_{i\neq j}\absolutevalue{a_{ij}}$ | term grows quickly with a denser $A$. This condition suggests that a more negative diagonal element is desirable, which ties in the concept of dissipation and its role in synchronization[31, 24, 32]. In addition, this condition fails to hold when $a_{jj}\geq 1$, due to the sum of absolute values always being nonnegative.
The second subcondition $\mu_{1}(A-I\circ A)<1$ appears more complicated, but is conceptually simple as well. Let the Hadamard product be represented by the operator $\circ$, so $I\circ A$ constructs a diagonal matrix with the diagonal elements of $A$. The resulting matrix $A-I\circ A$ is similar to $A$, but with its diagonal elements set to zero. This subcondition is a special case of (18), given by imposing the following constraint for every column:
$\sum_{i\neq j}\absolutevalue{a_{ij}}<1.$ (19)
Though (18) and (19) appear to overlap, these two subconditions disagree when the diagonal elements $a_{jj}$ are positive. Otherwise, if $a_{jj}$ are all nonpositive, then (19) implies (18).
The second condition for stability is $\max\{\mu_{\infty}(A),0\}<1$. Zero appears in this equation only to indicate that the leak term with matrix $C=I$, discussed in Section III.2, must be present to guarantee ESP and GS. Adhering to this constraint, the condition is effectively $\mu_{\infty}(A)<1$. This corresponds to the $\infty$-lognorm, which for an arbitrary matrix $M$ can be calculated by the following:
$\mu_{\infty}(M)=\max_{i}\left(m_{ii}+\sum_{j\neq i}\absolutevalue{m_{ij}}\right).$ (20)
The above condition is the row-wise version of (17), giving the corresponding restriction (18) when applied to the connectivity matrix $A$:
$\sum_{j\neq i}\absolutevalue{a_{ij}}<1-a_{ii}.$ (21)
@@ IV.1 UAP in Discrete Time
Taking the discrete time reservoir outlined in (1), we have established that $A:\norm{A}<1$ and $B$ can be arbitrarily chosen such that the ESP holds. The generation of trajectories is often called the listening phase and implies the existence of a unique time-invariant function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$ such that $x_{t}=\Phi(u_{t})$ as $t\to\infty$, but without an explicit representation of $\Phi$:
$x_{t+1}$ $=\sigma(Ax_{t}+Bu_{t}),$ (22)
$z_{t}$ $=Wx_{t}+b.$
@@
The constraint $\norm{A}<1$ allows for only one basin of attraction in the reservoir, resulting in the map $\Phi$ induced by the dynamics to also be unique. Aside from some transients, $u$ can be recovered from $x$ through the application of $\Phi^{-1}$, approximated by $W$. Further suppose that the approximation of $\Phi^{-1}$ by $W$ holds arbitrarily well, such that $u_{T}=\Phi^{-1}(x_{T})=Wx_{T}$ for a finite $T$. As such, the second line of (22) can be substituted into the first, resulting in an autonomous reservoir that is used for $t>T$. With the shorthand $\bar{A}=A+BW$ and $\bar{b}=Bb$, this newly formulated and autonomous reservoir is usually called the predicting phase:
@@
$=\sigma(\bar{A}x_{t}+\bar{b}),$
$\hat{u}_{t}$ $=Wx_{t}+b.$
It is surprising that a nonautonomous system like (22) can be converted into an autonomous system with a linear readout in the manner described in (23). Often overlooked is that the form of (23) admits the UAP in the feedforward sense of Cybenko [33], Hornik et al. [34], or Leshno et al. [35]. Approximation theorems for recurrent networks are not necessary.
Also overlooked is the relationship between contracting functions and the activation functions compatible with the UAP. Nonlinear contracting functions used as activation functions will always be compatible with the UAP. This is most succinctly shown with the results of Leshno et al. [35] which states that the UAP holds in networks of the form (23) if and only if the activation function is nonlinear and nonpolynomial.
The proof of nonlinear contracting functions being nonpolynomials is outlined as follows. If two distinct points on a smooth function cannot generate a slope greater than unity, then this function is globally contracting. Degree one and zero polynomials are linear, so they are automatically excluded from the set of nonlinear contracting functions. Polynomial functions of degree two or higher are nonlinear but will always manifest an unbounded slope eventually, so they cannot be contracting. Thus, the set of nonlinear contracting functions is mutually exclusive with the set of nonlinear polynomial functions. Therefore, all nonlinear contracting functions are nonpolynomials.
@@
The only invoked property of the activation function $\sigma$ thus far was nonlinear global contraction, which is a special case of Lipschitz continuous. Other commonly invoked properties of sigmoid functions – monotonicity and saturation – are not used in this paper, nor in Cybenko [33]. The UAP requires a large but finite number of nodes, hence the condition that $N\gg M$. In the spirit of Whitney (and consequently Takens when appropriate), almost every smooth map $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$ is an embedding when $N>2M$. See Theorem 2.2 of Sauer et al. [28] or Conjecture 2.3.4 of Hart et al. [17]. In the large $N$ limit, the approximation of $\Phi^{-1}(x_{t})\simeq Wx_{t}+b$ also holds arbitrarily well for ergodic systems [36].
@@ IV.2 UAP in continuous time
For continuous time reservoirs, let us first consider the Hopfield-like equations outlined in (7) with a leak current. We constrain $A$ such that $\norm{A}<1$ and $C=I$, but $B$ can be arbitrarily chosen such that the ESP holds. This implies the existence of some time-invariant function $\Phi:\mathbb{R}^{M}\to\mathbb{R}^{N}$ where $x(t)=\Phi(u(t))$ as $t\to\infty$ without explicit knowledge about the representation of $\Phi$:
$\dot{x}$ $=-x+\sigma(Ax+Bu),$ (24)
$\hat{u}$ $=Wx+b.$
Much like the discrete time case, trajectories are generated using the above dynamics, but an arbitrary time integration scheme is necessary. The listening and training phases are conducted as usual, where the weights $W$ and biases $b$ are solved for. The prediction phase is also done by converting the nonautonomous system into an autonomous one, via the closing of the input-output loop:
$\dot{x}$ $=-x+\sigma(\bar{A}x+\bar{b}),$ (25)
$\hat{u}$ $=Wx+b.$
@@
$\dot{x}$ $=\sigma(\bar{A}x+\bar{b}),$ (26)
$\hat{u}$ $=Wx+b.$
Matrix $A$ cannot yet be chosen such that the ESP is a priori guaranteed with the contraction approach following the arguments at the end of Section III. However, $A$ can be chosen such that the ESP occurs empirically for some input $u$ of interest. Assuming so, the reservoir trajectories are generated and the weights $W$ and bias $b$ are acquired by training. The resulting dynamics of the autonomous reservoir can be interpreted as a NODE with a linear output layer, but introduces another constraint in that the activation function are also invertible. There are multiple results [39, 40, 41, 42] that have established narrower forms of UAP for NODEs, which are inherited by autonomous reservoirs (26) with invertible activation functions. In particular, Zhang et al. [39] show that some augmented NODEs – augmented by imposing that the network has number of neurons at least twice the number of inputs – are capable of universally approximating homeomorphism of a topological space onto itself. Teshima et al. [40] show that NODEs are universal approximators of a large class of diffeomorphisms of a topological space onto itself.